# Distributed Multisensor ISAC

### v1

@@ VII Measured Example for Bistatic Joint Range-Doppler Estimation
Figure 24: Dual link ISAC demonstration setup. The radio nodes Rx1, Rx2 and Tx are stationary, the Target car was moving.
@@
Figure 26: Delay-Doppler spectrum with eToF and eDoppler paramter pair at Rx1.
It only appears as a small ripple. Even background subtraction cannot clearly reveal the target echo in the PDP, as seen in Figure 27. The PDPs in Figure 25 and Figure 27 have the same scaling in $\mathrm{dBm}\text{/}\mathrm{ns}$. The location of both LOS and target echo is indicated by the two colored markers. The two markers (turquoise-green and magenta) are at the same $\mathrm{dBm}\text{/}\mathrm{ns}$ level in both pictures. The difference is about $18\text{\,}\mathrm{dB}\text{/}$. The visibility of the target is enhanced but still bad.
However, the target reveals very well in the delay-Doppler distribution (also known as the delay-Doppler spectrum or the scattering function) which is calculated by the magnitude squared result of another set of iFFTs at every delay bin over a $50\text{\,}\mathrm{ms}\text{/}$ interval of slow time. The target echo is now very well separated by its Doppler shift while the static contribution (LOS and non-target related clutter) collapse at zero Doppler intersection in the delay-Doppler map. The target is clearly identified at $137\text{\,}\mathrm{ns}\text{/}$ eToF and $380\text{\,}\mathrm{Hz}\text{/}$ eDoppler (Rx1) and at $98\text{\,}\mathrm{ns}\text{/}$ eToF and $355\text{\,}\mathrm{Hz}\text{/}$ eDoppler (Rx2) in Figure 26 and Figure 28 respectively. From these two pairs measurements two eToF ellipse are drawn to identify the position of the car in 2D. The result is depicted in Figure 29.
@@
Unfortunately, the intersection is not unique and the GDoP would be high. Obviously, the position of the sensor nodes is not very appropriate. In an operational situation we would have to look for a better combination of nodes or to observe and track the position to enhance the result.
@@ VIII Outlook and Conclusion
MS ISAC has been shown to be the most promising and powerful approach to integrate radar sensing functionality into next generation mobile radio networks. While multiple radio nodes are readily available in cellular MU MIMO communication systems, MS ISAC also covers sidelink based mesh network approaches. However, use cases and sensing architectures (in terms of multi-sensor access and coordination, precoding, link adaptation, target parameter estimation and location tracking) will be very different if infrastructure only sensing, UL/DL sensing (with UE included as illuminator or sensor), or meshed sidelink sensing is pursued. Throughout this paper, we put a strong focus on the CPCL principle, first proposed in [ref1]. Not only does CPCL resemble and extend the classical passive radar paradigm, but it also appears to be pivotal in resolving bistatic synchronisation issues and enabling the proactive utilization of all mobile radio resources and functions for radar sensing and data processing.
With regard to the sensing function, MS ISAC constitutes a distributed MIMO radar. Multiple sensors brings the diversity advantage for more reliable target detection and better target recognition. We also gain the geometrical advantage for estimation and tracking of the 3D target state vector. However, MS ISAC has the additional benefit of reusing the ubiquitary existing, and constantly evolving resources of the mobile network. This gives us the performance gain in terms of coverage and resolution that can be achieved with the network densification and with other new 6G paradigms anyway (higher frequencies, wider bands, flexible radio access, precoding schemes, multiple antennas, including massive MIMO and RIS). In addition, heterogeneous radio interfaces and link adaptation tools, inherent data fusion and computation capabilities, and mobile sensor nodes that can be assigned a dedicated sensing mission, pave the way for a cognitive multifunction radar sensor network. We tried to unfold a comprehensive picture of MS ISAC. Because of the deep relation and the virtual equivalence of MS MIMO ISAC to MU MIMO communications we firmly belief that the first will play a similar role to ISAC as MU MIMO did to mobile communication.
However, we were unable to appreciate all the possible principles and advantages of MS ISAC on the same level. For example, the discussion on the large-scale use of AI to control resources and enable powerful estimation methods that can cope with interference and proactively exploit multipath propagation would deserve deeper discussion. In addition, many questions relating to multipath propagation were only addressed very briefly. For instance, the topics of targets bistatic reflectivity function measurement and modelling was not elaborated. This includes micro Doppler properties of targets that may be important for target detection, tracking and recognition [ref3, ref69, ref70, ref77, ref98].
@@
If we try to look into the future, we realize that the introduction of ISAC will be a longer process accompanied by technical and economic decisions. We will see these reflected in standardization processes. Since there are reservations about technical and regulatory issues, suspected impairments of communication performance, missing business models and even privacy concerns, it would be an exaggeration to expect that the road towards a perfect and comprehensive ISAC service would be fast track. However, there are also other examples (apart from pure voice telephony) in the long history of mobile communication that have finally caught on, albeit not in a straight line.
We think that infrastructure-only sensing, which does not directly include a user terminals under contract to the MNO in the sensing service cycle, could be a pilot example promising the fastest track for economic success. Reason is that ISAC is promising reasonable solutions for urgent needs in society, such as road and lower air traffic regulation, public security and safety, critical infrastructure protection, etc. Especially proprietary campus networks with dedicated applications and specific economic conditions could be a great place for ISAC use cases to grow.

### v2

@@ VII Measured Example for Bistatic Joint Range-Doppler Estimation
Figure 24: Dual link ISAC demonstration setup. The radio nodes Rx1, Rx2, and Tx are stationary; the target car was moving.
@@
Figure 26: Delay-Doppler spectrum with eToF and eDoppler parameter pair at Rx1.
It only appears as a small ripple. Even background subtraction cannot clearly reveal the target echo in the PDP, as seen in Figure 27. The PDPs in Figure 25 and Figure 27 have the same scaling in $\mathrm{dBm}\text{/}\mathrm{ns}$. The location of both LOS and the target echo is indicated by the two colored markers. The two markers (turquoise-green and magenta) are at the same $\mathrm{dBm}\text{/}\mathrm{ns}$ level in both pictures. The difference between them is about $18\text{\,}\mathrm{dB}\text{/}$. The visibility of the target is enhanced, but still bad.
However, the target reveals very well in the delay-Doppler distribution (also known as the delay-Doppler spectrum or the scattering function), which is calculated by the magnitude squared result of another set of FFTs at every delay bin over a $50\text{\,}\mathrm{ms}\text{/}$ interval of slow time according to the algorithm in Figure 15. The target echo is now very well separated by its Doppler shift, while the static contribution (LOS and non-target related clutter) collapses at the zero Doppler intersection in the delay-Doppler map. The target is clearly identified at $137\text{\,}\mathrm{ns}\text{/}$ eToF and $380\text{\,}\mathrm{Hz}\text{/}$ eDoppler (Rx1) and at $98\text{\,}\mathrm{ns}\text{/}$ eToF and $355\text{\,}\mathrm{Hz}\text{/}$ eDoppler (Rx2) in Figure 26 and Figure 28 respectively. From these two measurements, two eToF ellipses are drawn to identify the position of the car in 2D. The result is depicted in Figure 29. Unfortunately, the intersection is not unique, and the GDoP would be high. This tells us that the positions of the sensor nodes in our example was not very appropriate. In an operational situation, we would have to look for a better combination of nodes or observe and track the position to enhance the result.
@@ VIII Outlook and Conclusion
MS-ISAC has been shown to be the most promising and powerful approach to integrate radar sensing functionality into next generation mobile radio networks. While multiple radio nodes are readily available in cellular MU MIMO communication systems, MS-ISAC also covers sidelink based mesh network approaches. However, use cases and sensing architectures (in terms of multi-sensor access and coordination, precoding, link adaptation, target parameter estimation and location tracking) will be very different if infrastructure only sensing, UL/DL sensing (with UE included as illuminator or sensor), or meshed sidelink sensing is pursued. Throughout this paper, we put a strong focus on the CPCL principle, first proposed in [ref1]. Not only does CPCL resemble and extend the classical passive radar paradigm, but it also appears to be pivotal in resolving bistatic synchronisation issues and enabling the proactive utilization of all mobile radio resources and functions for radar sensing and data processing.
With regard to the sensing function, MS-ISAC constitutes a distributed MIMO radar. Multiple sensors bring the diversity advantage for more reliable target detection and better target recognition. We also gain the geometrical advantage for estimation and tracking of the 3D target state vector. MS-ISAC has the additional benefit of reusing the ubiquitary existing, and constantly evolving resources of the mobile network. This gives us the performance gain in terms of coverage and resolution that can be achieved with the network densification and with other new 6G paradigms anyway (higher frequencies, wider bands, flexible radio access, precoding schemes, multiple antennas, including massive MIMO and RIS). In addition, heterogeneous radio interfaces and link adaptation tools, inherent data fusion and computation capabilities, and mobile sensor nodes that can be assigned a dedicated sensing mission, pave the way for a cognitive multifunction radar sensor network. We tried to unfold a comprehensive picture of MS-ISAC. Because of the deep relation and the virtual equivalence of MS MIMO ISAC to MU MIMO communications we firmly belief that the first will play a similar role to ISAC as MU MIMO did to mobile communication.
However, we were unable to appreciate all the possible principles and advantages of MS-ISAC on the same level. For example, the discussion on the large-scale use of AI to control resources and enable powerful estimation methods that can cope with interference and proactively exploit multipath propagation would deserve deeper discussion. In addition, many questions relating to multipath propagation were only addressed very briefly. For instance, the topics of targets’ bistatic reflectivity function measurements and modeling would need deeper discssion. This includes micro Doppler properties of targets that may be important for target detection, tracking and recognition [ref3, ref69, ref70, ref77, ref98].
@@
Cosidering future developments, we realize that the introduction of ISAC will be a longer process accompanied by technical and economic decisions. We will see these reflected in standardization processes. Since there are reservations about technical and regulatory issues, suspected impairments of communication performance, missing business models and even privacy concerns, it would be an exaggeration to expect that the road towards a perfect and comprehensive ISAC service would be a fast track. However, there are also other examples (apart from pure voice telephony) in the long history of mobile communication that have finally caught on, albeit not in a straight line.
We think that infrastructure-only sensing, which does not directly include a user terminal under contract to the MNO in the sensing service cycle, could be a pilot example promising the fastest route to economic success. Reason is that ISAC is promising and reasonable solutions for urgent needs in society, such as road and lower air traffic regulation, public security and safety, critical infrastructure protection, etc. Especially proprietary campus networks with dedicated applications and specific economic conditions could be a great place for ISAC use cases to grow. For public networks, operator models of ISAC based surveillance services to support public safety issues could be developed similar to other Mission Critical Services (MCX) in 5G.